OpenStudy (anonymous):

f(x)=x^3-7x+6 find the rational zeros.

OpenStudy (anonymous):

Well by inspection you can see that 1 is a solution 1-7+6=0 so: \[x^3-7x+6=(x-1)(x^2+x-6)=(x-1)(x+3)(x-2)=0 \] \[\implies x=1,2,-3\]